OpenStudy (anonymous):

Write the solution in lowest terms. 1/3 - (-3/4) = A. 4/12 B. C. 1/3

OpenStudy (anonymous):

1/3-(-3/4)=1/3+3/4=4/12+9/12=13/12

OpenStudy (anonymous):

Though it doesn't seem to be one of the solutions... What is B?

OpenStudy (anonymous):

1 1/12

OpenStudy (anonymous):

Ah, well in that case, 13/12 is also equal to 12/12+1/12, which is equal to 1 and 1/12; so B is the answer

OpenStudy (anonymous):

okay thank you